# Emergence of Fixational and Saccadic Movements in a Multi-Level Recurrent Attention Model for Vision

### v1

@@ Abstract
Keywords:
Image classification Active Vision Recurrent Neural Networks Hard Attention Biologically Inspired Computation
@@ 1 Introduction
Human vision actively interacts with the environment by eye movement, characterized by recurrent fixations on specific spatial locations and saccadic movements. Rather than passively receiving uniform visual input, human eyes deploy focused glimpses of high-resolution processing in the centered fovea vision, complemented by peripheral. This gazed information integrates over time, allowing humans to understand and explore relevant scene regions in a sequential, yet highly flexible manner. Seminal work by Yarbus demonstrates that human eye scanpaths, formed by sequences of fixations and saccades, reflect not only the viewer’s focus but also mental activities such as the nature of the task [35]. Although factors such as the ongoing tasks and the expert knowledge of humans have an impact on the scan paths, human eye movement shares similarities in the fixational and saccadic patterns [28, 17].
@@
Studies in neuroscience reveal the cortical pathway of human visual information processing. Human visual processing includes two distinct pathways from the retina, one being the geniculostriate pathway, that information going through the lateral geniculate nucleus (LGN) to the primary visual cortex, and another, the eye movement pathway, directly going to the superior colliculus (SC) along the optic nerve, projecting to the pons paramedian pontine reticular formation that sends control signals to oculomotor nuclei controlling the eye movement [22]. The processed information in the primary visual cortex, posterior parietal cortex, and frontal eye fields also projects back to the SC for eye movement control. It was shown by lesions of the visual cortex causing a loss of peripheral vision while the center vision remains unaffected. This suggests that the center vision relies mainly on the visual information directly received by the retina to the SC for eye movement control, while the geniculostriate pathway processes the center vision and peripheral vision and provides feedback for indirect eye movement control.
@@
Recent advances in machine learning and computational methods of visual modeling focus on the whole image processing. The Convolutional Neural Network (CNN) model has shown high accuracy on visual tasks and similarity in the visual information processing in the visual cortex [24]. However, such computation usually has a high computation cost and a black-box model that is insufficient in interpreting its prediction. In addressing these problems, attention models in machine learning seek to replicate the nature of foveal vision and attention mechanisms. While soft attention employs a standard CNN to distribute the attention map over the entire input, hard attention models employ a learnable policy to select small sub-regions or “glimpses” and process them with higher resolution. The hard attention strategy can reduce computational demand, improve interpretability, and learn a dynamic human-like gaze behavior. Even though hard attention is harder to train than soft attention because the cut glimpses are non-differentiable. Seminal work by Mnih et al. proposed the Recurrent Model of Visual Attention (RAM), showing one can learn hard attention through the Reinforcement Learning (RL) method [26]. They train a recurrent network to sequentially choose glimpse locations and to perform tasks such as classification from partial observations with the REINFORCE algorithm [31]. The subsequent variations, such as the Deep Recurrent Attention Model (DRAM), are proposed to solve dedicated visual tasks such as multi-object recognition.
@@ 2.1 Recurrent Models of Visual Attention
Visual attention has long stood as a cornerstone topic in computer vision research. Studies from neuroscience indicate that signals guided by prior knowledge interact with those influenced by raw input, creating a competitive process to select relevant elements in a scene [9, 3, 2, 10]. Early efforts in this area introduced saliency-based approaches, usually determined by local low-level feature contrast [21]. Incorporating feedback neural networks, recurrent networks, and reinforcement learning into attention-based networks has led to noteworthy gains across a variety of visual tasks [5, 14, 27, 4].
Recent methods evolved to allocate processing resources more efficiently. For instance, hard attention strategies limit computation to specific parts of an image, reducing computational demands [23, 33]. The RAM then broadened this idea, using sequential glimpses to examine visual content while preserving interpretability [26]. Following RAM, the DRAM was proposed, which used multiple glimpses in deeper recurrent structures, showing improvements on multi-digit recognition tasks such as the Street View House Numbers dataset [1]. More recent variants of RAM extended the hard attention to dynamically adjusting the stop time, modeling the saccadic behavior by restraining the model from visited locations, and a biological approach using predictive coding [25, 13, 29].
@@ 2.2 Hierarchical and Multi-Scale Extensions
Hierarchical strategies have gained attention due to the layered nature of tasks and the need to accommodate targets at varying levels of detail. One early demonstration of this principle introduced a tiered feature extraction pipeline, successfully capturing both low-level cues and higher-level semantics for object recognition and segmentation [16]. Meanwhile, modeling how people look around in the context of goals and sub-goals also highlights the importance of structured, task-driven control of eye movements [18].
Within deep learning, sequential models have embraced multi-scale designs that split complex input into progressively more abstract representations. For instance, a layered recurrent network has been shown to improve skeletal action recognition by separately capturing local movements before integrating them globally [11]. A related design relies on an attention mechanism stacked at different levels—focusing first on words, then on sentences—to boost performance in document classification [34]. Extensions of this theme push further into multi-resolution contexts, dealing effectively with intricate temporal or spatial patterns [6, 7].
Beyond supervised approaches, reinforcement learning has also taken advantage of hierarchical structures. Subdividing a large goal into manageable chunks can accelerate training and produce more robust behaviors [19]. In general, these hierarchical and multi-scale frameworks have shown notable improvements in tackling visual challenges, reflecting the real-world complexity of tasks and their multi-level requirements. Therefore, a hierarchical RAM model aligns more closely with the way humans move their eyes while also leveraging the tiered nature of visual and task demands, ultimately enhancing performance in our approach.
@@ 3.2 MRAM
We introduce two recurrent layers, with the Long-Short Term Memory (LSTM) model [20] serving as the RNN at each layer, denoted $h^{1}_{t}$ and $h^{2}_{t}$. At each time step $t$:
@@
    Hierarchical factorisation of POMDP. In MRAM, we model visual search as a partially–observable MDP with latent state $s_{t}$, observation $o_{t}$, gaze action $g_{t}$, and terminal classification $y$. In a flat RAM, a single RNN must learn the full joint policy $\pi_{\theta}(g_{1:T},y\!\mid\!o_{1:T})$, entangling short–horizon saccade control with long–horizon recognition. In DRAM, the lower RNN learn long–horizon recognition, while the higher RNN learn the short–horizon saccade control. The mismatch in time-scale leads to compromised performance on both gaze action and classification. We instead factorise the policy into a two-layer hierarchy (SC–layer ${\pi_{L}}$, VC–layer ${\pi_{H}}$):
@@
    where $\pi_{L}$ evolves on a fast time–scale (saccadic glimpse control, superior colliculus) and $\pi_{H}$ on a slow time–scale (object recognition, visual cortex). The factorisation is lossless yet separates exploration ($\pi_{L}$) from exploitation ($\pi_{H}$), reducing credit-assignment span like typical hierarchical reinforcement learning method did.
@@ 3.3 Hybrid Baseline Estimation
To reduce variance in policy gradient training, we applied a baseline function $b_{t}=\mathop{\mathbb{E}}_{\pi}[R_{t}]$ that approximates the expected return [30]. Unlike standard baselines in RAM or DRAM, which often use a single hidden state, our hybrid baseline fuses the hidden states from both layers, resembling the feedback projection from the primary visual cortex to SC:
@@
[…]
@@ 3.4 Training
We follow a policy gradient approach for learning the location policy akin to the same REINFORCE algorithm applied in RAM [26, 31]. The location policy for the lower layer is a stochastic component, which samples the glimpse location $\ell_{t+1}$ from a distribution parameterized by $h^{1}_{t}$. The action module of the higher layer is learned in a supervised approach with the classification outputs $y_{t}$ and the ground-truth label. The overall hybrid loss combines:
@@ 4.1 MNIST
In this task, we evaluate the model in same standard dataset used in RAM, MNIST datasets. The MNIST dataset is handwritten images ($28\times 28$) with digital numbers ranging from 0 to 9 [8]. We evaluate the contribution of each module to MRAM with ablation on the CNN and baseline. We record the performance of different models and the effectiveness of individual modules such as total trainable parameters and average training time of a single epoch. We compare our MRAM model to RAM and DRAM under identical training conditions, with a small LeNet-5 as CNN baseline, reporting test set accuracy in Tables 1.
@@
While RAM is reported the best performance at the 7 glimpse steps, we tested the glimpses to 7 and 10 for comparing short and long glimpse. The glimpsed image size is set to be fixed at $8\times 8$ same as reported in RAM[26]. The number of scale is set to 1. MRAM achieves a state-of-the-art hard attention accuracy of 99.2% at 10 glimpse steps, outperforming all baseline models. We also evaluated the contribution of the context, a full image global information extracted by a CNN and passed to the initial state of the higher layer in DRAM as presented in DRAM[1]. The DRAM relies on the context information, and removing the context module leads to a 0.47% accuracy loss. To justify the contribution of the hybrid baseline method in MRAM, we tested the MRAM train with a single baseline method, the same as the RAM and DRAM models, and the DRAM to train with the hybrid baseline method in the 10 glimpses experiment setting. The MRAM loses 0.25% accuracy, while DRAM loses 0.06% accuracy. This suggests the effectiveness of the hybrid baseline method inspired by the feedback projection from the visual cortex to the SC fit only on the MRAM model. Longer glimpse step leads to longer training cost. The cost per glimpse for of RAM/MRAM in training is about 11s/epoch, which is smaller than DRAM (13s/epoch) and CNN, while the inference time doesn’t vary much(Tab.1).
@@ 4.2 FashionMNIST
The FashionMNIST dataset is MNIST-like grayscale images ($28\times 28$) with pictures of common clothes categories such as dress and sneaker [32]. To further explore the performance on image classification and the learned glimpse policy, we record the parameters, inference time per image, accuracy of different models and the predicted locations on the dataset. The baseline, all parameters are the same with MNIST experiment, except we increased the longer glimpse step from 10 to 12 steps for larger gap between short and long glimpses.
@@ 4.3 Self-Emergence of Fixations and Saccades in Category Classification
To answer the question of how MRAM outperformed RAM and whether a human-like gaze policy is learned, we examine the sequentially predicted glimpse locations during the classification task in FashionMNIST. Figure 2 illustrates examples of the predicted fixation locations for different item categories in FashionMNIST. Notably, MRAM often employs repeated fixations around distinctive features of an object and occasionally executes longer saccades to explore new regions. In contrast, the RAM model and DRAM model learn a distinct glimpse policy: one learns to continuously place glimpses in new locations like saccades, while the other learns to focus on a relatively small region like fixation. The concentrated glimpse policy suggested DRAM failed to learn the proper gaze policy because of the overwhelmingly powerful global context provided by CNN. RAM typically reach peak performance at a specific number of glimpses, after which additional steps may provide diminishing returns or even degrade performance. This observation has motivated variants such as Dynamic RAM, which adaptively adjust the number of glimpses [25]. It can be explained by our visualization result that RAM is always predicting gaze like saccade, and a manual stop is necessary before the saccade moved away from target. In contrast, our MRAM framework can sustain or improve its classification accuracy as the number of glimpses increases, leveraging its hierarchical structure to incorporate both fixations and broader saccadic shifts. In general, MRAM’s hierarchical design, inspired by the human brain, enables it to maintain robustness by refining local information across glimpses while also making larger saccades when needed.
@@
To quantitatively explain the visualization of the learned policy of each model, we showed the probability density of fixation durations and jump distance with the kernel distance estimate on the predicted glimpse locations. As the glimpse windows are $8\times 8$, it’s natural to define a fixation by the distance of two glimpses lower than 8. We defined the fixation with a more strict threshold distance of 6. The probability density of fixation durations and jump distances is shown in Figure 3. MRAM learned a glimpse policy with short fixation durations but longer jump distances compared to DRAM. This emerged pattern echoes the findings of Yarbus [35] and the fixations duration and saccade distance data of human vision in [15], in which humans combine long-time fixations and sporadic long-range gaze shifts. Categories such as dresses and shirts, which span a significant vertical portion of the image, benefit especially from multiple fixations capturing various sections of the item’s silhouette. Our method thus learns to focus on salient or discriminative regions repeatedly while also retaining the flexibility to jump to new areas when needed, ultimately leading to improved recognition performance.
@@ 4.4 FER2013
To evaluate our model on real-world image, we tested our models and baseline models on the Facial Expression Recognition (FER) 2013 dateset with ($48\times 48$) images of human face[12]. As we tested that longer glimpse is beneficial for all models in FashionMNIST, we fix the glimpse steps to 12, and the glimpsed image size is set to be fixed at $8\times 8$. In the 2 scale setting, the peripheral vision of $16\times 16$ image patch will be down-sampled to $8\times 8$, and concatenated as the image input to the RAM-based models.

### v2

@@ 1 Introduction
Human vision actively interacts with the environment by eye movement, characterized by recurrent fixations on specific spatial locations and saccadic movements. Rather than passively receiving uniform visual input, human eyes deploy focused glimpses of high-resolution processing in the centered fovea vision, complemented by peripheral. This gazed information integrates over time, allowing humans to understand and explore relevant scene regions in a sequential, yet highly flexible manner. Seminal work by Yarbus demonstrates that human eye scanpaths, formed by sequences of fixations and saccades, reflect not only the viewer’s focus but also mental activities such as the nature of the task [Yarbus]. Although factors such as the ongoing tasks and the expert knowledge of humans have an impact on the scan paths, human eye movement shares similarities in the fixational and saccadic patterns [Eye1, Eye2].
@@
Studies in neuroscience reveal the cortical pathway of human visual information processing. Human visual processing includes two distinct pathways from the retina, one being the geniculostriate pathway, that information going through the lateral geniculate nucleus (LGN) to the primary visual cortex, and another, the eye movement pathway, directly going to the superior colliculus (SC) along the optic nerve, projecting to the pons paramedian pontine reticular formation that sends control signals to oculomotor nuclei controlling the eye movement [Kandel]. The processed information in the primary visual cortex, posterior parietal cortex, and frontal eye fields also projects back to the SC for eye movement control. It was shown by lesions of the visual cortex causing a loss of peripheral vision while the center vision remains unaffected. This suggests that the center vision relies mainly on the visual information directly received by the retina to the SC for eye movement control, while the geniculostriate pathway processes the center vision and peripheral vision and provides feedback for indirect eye movement control.
@@
Recent advances in machine learning and computational methods of visual modeling focus on the whole image processing. The Convolutional Neural Network (CNN) model has shown high accuracy on visual tasks and similarity in the visual information processing in the visual cortex [CNN]. However, such computation usually has a high computation cost and a black-box model that is insufficient in interpreting its prediction. In addressing these problems, attention models in machine learning seek to replicate the nature of foveal vision and attention mechanisms. While soft attention employs a standard CNN to distribute the attention map over the entire input, hard attention models employ a learnable policy to select small sub-regions or “glimpses” and process them with higher resolution. The hard attention strategy can reduce computational demand, improve interpretability, and learn a dynamic human-like gaze behavior. Even though hard attention is harder to train than soft attention because the cut glimpses are non-differentiable. Seminal work by Mnih et al. proposed the Recurrent Model of Visual Attention (RAM), showing one can learn hard attention through the Reinforcement Learning (RL) method [RAM]. They train a recurrent network to sequentially choose glimpse locations and to perform tasks such as classification from partial observations with the REINFORCE algorithm [REINFORCE]. The subsequent variations, such as the Deep Recurrent Attention Model (DRAM), are proposed to solve dedicated visual tasks such as multi-object recognition.
@@ 2.1 Recurrent Models of Visual Attention
Visual attention has long stood as a cornerstone topic in computer vision research. Studies from neuroscience indicate that signals guided by prior knowledge interact with those influenced by raw input, creating a competitive process to select relevant elements in a scene [attention1, attention2, biased1, biased2]. Early efforts in this area introduced saliency-based approaches, usually determined by local low-level feature contrast [saliency]. Incorporating feedback neural networks, recurrent networks, and reinforcement learning into attention-based networks has led to noteworthy gains across a variety of visual tasks [attention3, attention4, attention5, RLattention].
Recent methods evolved to allocate processing resources more efficiently. For instance, hard attention strategies limit computation to specific parts of an image, reducing computational demands [fovea, Xu]. The RAM then broadened this idea, using sequential glimpses to examine visual content while preserving interpretability [RAM]. Following RAM, the DRAM was proposed, which used multiple glimpses in deeper recurrent structures, showing improvements on multi-digit recognition tasks such as the Street View House Numbers dataset [DRAM]. More recent variants of RAM extended the hard attention to dynamically adjusting the stop time, modeling the saccadic behavior by restraining the model from visited locations, and a biological approach using predictive coding [DyRAM, saccader, PCRAM].
@@ 2.2 Hierarchical and Multi-Scale Extensions
Hierarchical strategies have gained attention due to the layered nature of tasks and the need to accommodate targets at varying levels of detail. One early demonstration of this principle introduced a tiered feature extraction pipeline, successfully capturing both low-level cues and higher-level semantics for object recognition and segmentation [hierarchy1]. Meanwhile, modeling how people look around in the context of goals and sub-goals also highlights the importance of structured, task-driven control of eye movements [hierarchy2].
Within deep learning, sequential models have embraced multi-scale designs that split complex input into progressively more abstract representations. For instance, a layered recurrent network has been shown to improve skeletal action recognition by separately capturing local movements before integrating them globally [HRNN1]. A related design relies on an attention mechanism stacked at different levels—focusing first on words, then on sentences—to boost performance in document classification [HAN]. Extensions of this theme push further into multi-resolution contexts, dealing effectively with intricate temporal or spatial patterns [HRNN2, HRNN3].
Beyond supervised approaches, reinforcement learning has also taken advantage of hierarchical structures. Subdividing a large goal into manageable chunks can accelerate training and produce more robust behaviors [HRL]. In general, these hierarchical and multi-scale frameworks have shown notable improvements in tackling visual challenges, reflecting the real-world complexity of tasks and their multi-level requirements. Therefore, a hierarchical RAM model aligns more closely with the way humans move their eyes while also leveraging the tiered nature of visual and task demands, ultimately enhancing performance in our approach.
@@ 3.2 MRAM
We introduce two recurrent layers, with the Long-Short Term Memory (LSTM) model [LSTM] serving as the RNN at each layer, denoted $h^{1}_{t}$ and $h^{2}_{t}$. At each time step $t$:
@@
    Hierarchical factorisation of POMDP. In MRAM, we model visual search as a partially–observable MDP with latent state $s_{t}$, observation $o_{t}$, gaze action $g_{t}$, and terminal classification $y$. In a flat RAM, a single RNN must learn the full joint policy $\pi_{\theta}(g_{1:T},y\!\mid\!o_{1:T})$, entangling short–horizon saccade control with long–horizon recognition. In DRAM, the lower RNN learn long–horizon recognition, while the higher RNN learn the short–horizon saccade control. The mismatch in time-scale leads to compromised performance on both gaze action and classification. We instead factorise the policy into a two‑layer hierarchy (SC–layer $\pi_{L}$}, VC–layer $\pi_{H}$}):
@@
    where $\pi_{L}$ evolves on a fast time–scale (saccadic glimpse control, superior colliculus) and $\pi_{H}$ on a slow time–scale (object recognition, visual cortex). The factorisation is lossless yet separates exploration ($\pi_{L}$) from exploitation ($\pi_{H}$), reducing credit‑assignment span like typical hierarchical reinforcement learning method did.
@@ 3.3 Hybrid Baseline Estimation
To reduce variance in policy gradient training, we applied a baseline function $b_{t}=\mathop{\mathbb{E}}_{\pi}[R_{t}]$ that approximates the expected return [baseline]. Unlike standard baselines in RAM or DRAM, which often use a single hidden state, our hybrid baseline fuses the hidden states from both layers, resembling the feedback projection from the primary visual cortex to SC:
@@
where $\|$” denotes concatenation and $f_{\mathrm{b}}$ is a small multi-layer perceptron producing the scalar baseline. This design leverages both low-level and higher-level cues, providing a more stable estimate for the reward expectation.
@@ 3.4 Training
We follow a policy gradient approach for learning the location policy akin to the same REINFORCE algorithm applied in RAM [RAM, REINFORCE]. The location policy for the lower layer is a stochastic component, which samples the glimpse location $\ell_{t+1}$ from a distribution parameterized by $h^{1}_{t}$. The action module of the higher layer is learned in a supervised approach with the classification outputs $y_{t}$ and the ground-truth label. The overall hybrid loss combines:
@@ 4.1 MNIST
In this task, we evaluate the model in same standard dataset used in RAM, MNIST datasets. The MNIST dataset is handwritten images ($28\times 28$) with digital numbers ranging from 0 to 9 [MNIST]. We evaluate the contribution of each module to MRAM with ablation on the CNN and baseline. We record the performance of different models and the effectiveness of individual modules such as total trainable parameters and average training time of a single epoch. We compare our MRAM model to RAM and DRAM under identical training conditions, with a small LeNet-5 as CNN baseline, reporting test set accuracy in Tables 1.
@@
While RAM is reported the best performance at the 7 glimpse steps, we tested the glimpses to 7 and 10 for comparing short and long glimpse. The glimpsed image size is set to be fixed at $8\times 8$ same as reported in RAM[RAM]. The number of scale is set to 1. MRAM achieves a state-of-the-art hard attention accuracy of 99.2% at 10 glimpse steps, outperforming all baseline models. We also evaluated the contribution of the context, a full image global information extracted by a CNN and passed to the initial state of the higher layer in DRAM as presented in DRAM[DRAM]. The DRAM relies on the context information, and removing the context module leads to a 0.47% accuracy loss. To justify the contribution of the hybrid baseline method in MRAM, we tested the MRAM train with a single baseline method, the same as the RAM and DRAM models, and the DRAM to train with the hybrid baseline method in the 10 glimpses experiment setting. The MRAM loses 0.25% accuracy, while DRAM loses 0.06% accuracy. This suggests the effectiveness of the hybrid baseline method inspired by the feedback projection from the visual cortex to the SC fit only on the MRAM model. Longer glimpse step leads to longer training cost. The cost per glimpse for of RAM/MRAM in training is about 11s/epoch, which is smaller than DRAM (13s/epoch) and CNN, while the inference time doesn’t vary much(Tab.1).
@@ 4.2 FashionMNIST
The FashionMNIST dataset is MNIST-like grayscale images ($28\times 28$) with pictures of common clothes categories such as dress and sneaker [FashionMNIST]. To further explore the performance on image classification and the learned glimpse policy, we record the parameters, inference time per image, accuracy of different models and the predicted locations on the dataset. The baseline, all parameters are the same with MNIST experiment, except we increased the longer glimpse step from 10 to 12 steps for larger gap between short and long glimpses.
@@ 4.3 Self-Emergence of Fixations and Saccades in Category Classification
To answer the question of how MRAM outperformed RAM and whether a human-like gaze policy is learned, we examine the sequentially predicted glimpse locations during the classification task in FashionMNIST. Figure 2 illustrates examples of the predicted fixation locations for different item categories in FashionMNIST. Notably, MRAM often employs repeated fixations around distinctive features of an object and occasionally executes longer saccades to explore new regions. In contrast, the RAM model and DRAM model learn a distinct glimpse policy: one learns to continuously place glimpses in new locations like saccades, while the other learns to focus on a relatively small region like fixation. The concentrated glimpse policy suggested DRAM failed to learn the proper gaze policy because of the overwhelmingly powerful global context provided by CNN. RAM typically reach peak performance at a specific number of glimpses, after which additional steps may provide diminishing returns or even degrade performance. This observation has motivated variants such as Dynamic RAM, which adaptively adjust the number of glimpses [DyRAM]. It can be explained by our visualization result that RAM is always predicting gaze like saccade, and a manual stop is necessary before the saccade moved away from target. In contrast, our MRAM framework can sustain or improve its classification accuracy as the number of glimpses increases, leveraging its hierarchical structure to incorporate both fixations and broader saccadic shifts. In general, MRAM’s hierarchical design, inspired by the human brain, enables it to maintain robustness by refining local information across glimpses while also making larger saccades when needed.
@@
To quantitatively explain the visualization of the learned policy of each model, we showed the probability density of fixation durations and jump distance with the kernel distance estimate on the predicted glimpse locations. As the glimpse windows are $8\times 8$, it’s natural to define a fixation by the distance of two glimpses lower than 8. We defined the fixation with a more strict threshold distance of 6. The probability density of fixation durations and jump distances is shown in Figure 3. MRAM learned a glimpse policy with short fixation durations but longer jump distances compared to DRAM. This emerged pattern echoes the findings of Yarbus [Yarbus] and the fixations duration and saccade distance data of human vision in [Saccadepattern1], in which humans combine long-time fixations and sporadic long-range gaze shifts. Intriguingly, we reproduced the DRAM-like behavior on MRAM, when the locator uses only output from the higher recurrent layer during test(MRAM r2 only). This ablation support that human-like eye movement comes from our hierarchical structure decoupling 2 RNN layers for 2 distinct functions. Categories such as dresses and shirts, which span a significant vertical portion of the image, benefit especially from multiple fixations capturing various sections of the item’s silhouette. Our method thus learns to focus on salient or discriminative regions repeatedly while also retaining the flexibility to jump to new areas when needed, ultimately leading to improved recognition performance.
@@ 4.4 FER2013
To evaluate our model on real-world image, we tested our models and baseline models on the Facial Expression Recognition (FER) 2013 dateset with ($48\times 48$) images of human face[FER]. As we tested that longer glimpse is beneficial for all models in FashionMNIST, we fix the glimpse steps to 12, and the glimpsed image size is set to be fixed at $8\times 8$. In the 2 scale setting, the peripheral vision of $16\times 16$ image patch will be down-sampled to $8\times 8$, and concatenated as the image input to the RAM-based models.